# EventFlow: Forecasting Temporal Point Processes with Flow Matching

### v2

@@ 4 EventFlow
Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages. First, EventFlow directly models the joint distribution over event times, thereby avoiding autoregression entirely. Second, our model is likelihood-free at training time, avoiding the Monte Carlo estimates needed to estimate the likelihood in Equation (2). Third, sampling from our model amounts to solving an ordinary differential equation. This is straightforward to implement and parallelize, allowing us to avoid the difficulties of thinning-based approaches. In fact, EventFlow can achieve competitive performance using only a single forward pass at sampling time. We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model. We begin below by focusing on the unconditional setting, followed by an extension for conditional generation.
@@
We first introduce some necessary background and notation. Let $\mathcal{T}=[0,T]$ be a finite length interval. The set $\Gamma$ denotes the configuration space of $\mathcal{T}$ (Albeverio et al., 1998), i.e., the set of all finite counting measures on the set $[0,T]$. A point $\gamma\in\Gamma$ corresponds to a measure of the form $\gamma=\sum_{k=1}^{n}\delta[t^{k}]$, i.e., a finite collection of Dirac deltas located at event times $t^{k}\in\mathcal{T}$. A temporal point process (TPP) on $\mathcal{T}$ is a probability distribution $\mu$ over the configuration space $\Gamma$. Informally, $\mu$ represents a distribution over finite sequences $\gamma\in\Gamma$ whose events are in $\mathcal{T}$. We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the sequence $\gamma$ [^1] This can be thought of in terms of the counting process, i.e., $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$. While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg, 2017).
Under mild assumptions, a TPP $\mu$ can be fully characterized (Daley & Vere-Jones, 2003, Prop. 5.3.II) by a probability distribution which specifies the number of events and a collection of joint densities corresponding to the event times themselves. In a slight abuse of notation, we will write $\mu(n)$ for the corresponding distribution over event counts, and $\{\mu^{n}(t^{1},\dots,t^{n})\}_{n=1}^{\infty}$ for the collection of joint distributions. In other words, for any given $n\in\mathbb{Z}_{\geq 0}$, the probability of observing $n$ events in the interval $\mathcal{T}$ is $\mu(n)$, and $\mu^{n}(t^{1},\dots,t^{n})$ describes the corresponding joint distribution of event times. We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that $t^{1}<t^{2}<\dots<t^{n}$.
Let $\mu_{1}$ represent the data distribution and $\mu_{0}$ represent a reference distribution. That is, both $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ are TPP distributions. To construct our model, we will define a path of TPPs $\eta_{s}\in\mathbb{P}(\Gamma)$ which approximately interpolates from our reference distribution to our data distribution. Throughout, we use $s\in[0,1]$ to denote a flow time and $t\in[0,T]$ to denote an event time. These two time axes are in a sense orthogonal to one another (see Figure 1).
@@
Our first step is to define a useful notion of couplings, allowing us to pair event sequences drawn from $\mu_{0}$ with those drawn from $\mu_{1}$. A coupling between two TPPs $\mu,\nu\in\mathbb{P}(\Gamma)$ is a joint probability measure $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu$ and $\nu$. We say that the coupling $\rho$ is balanced if draws $(\gamma_{0},\gamma_{1})\sim\rho$ are such that $N(\gamma_{0})=N(\gamma_{1})$ almost surely. In other words, balanced couplings only pair event sequences with equal numbers of events. While we will later see how to interpolate between any two given event sequences, this coupling will allow us to decide which sequences to interpolate. In particular, a balanced coupling will allow us to pair sequences such that they always have the same number of events, allowing us to avoid the addition or deletion of events during both training and sampling. We will use $\Pi_{b}(\mu,\nu)$ to denote the set of balanced couplings. The following proposition shows $\Pi_{b}(\mu,\nu)$ is nonempty if and only if the event count distributions of $\mu$ and $\nu$ are equal, placing a structural constraint on the suitable choices of $\mu_{0}$.
@@
Let $\mu,\nu\in\mathbb{P}(\Gamma)$ be two TPPs. The set of balanced couplings $\Pi_{b}(\mu,\nu)$ is nonempty if and only if $\mu(n)=\nu(n)$ have the same distribution over event counts.
We provide a proof in Appendix B. In practice, we follow a simple strategy for choosing both the reference TPP $\mu_{0}$ and the coupling $\rho$. Suppose $q$ is any given density on our state space $\mathcal{T}$, e.g., a uniform distribution. We take $\mu_{0}$ to be a mixed binomial process (Kallenberg, 2017, Ch. 3) whose event count distribution is given by that of the data $\mu_{1}(n)$ and joint event distributions given by independent products of $q$ (up to sorting). That is, to sample from $\mu_{0}$, one can simply sample $n\sim\mu_{1}(n)$ from the empirical event count distribution followed by sampling and sorting $n$ i.i.d. points $t^{k}\sim q$. To draw a sample from our coupling $\rho$, we first sample a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a draw $\gamma_{0}\sim\mu_{0}$.
@@
We now proceed to construct our interpolant $\eta_{s}\in\mathbb{P}(\Gamma)$. We will construct this path of TPPs via a local procedure which we then marginalize over a given balanced coupling. Here, we extend flow matching techniques (Lipman et al., 2023; Tong et al., 2024) to design our sequence-level interpolants, but we emphasize that this is only possible under a balanced coupling as the number of events is fixed. To that end, let $\rho$ be any balanced coupling of the reference measure $\mu_{0}$ and the data measure $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling. As $\rho$ is balanced, we have $\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]$ and $\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ are both a collection of $n$ events. As TPPs are fully characterized by their joint distributions over event times, we will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$. First, we the define measure $\gamma_{s}^{z}\in\Gamma$ via
@@
where we use the superscript $z$ to denote the dependence on the pair $z=(\gamma_{0},\gamma_{1})$. In other words, $\gamma_{s}^{z}$ linearly interpolates each corresponding event in $\gamma_{0}$ and $\gamma_{1}$. This defines a path $(\gamma_{s}^{z})_{s=0}^{1}$ in the configuration space $\Gamma$ which evolves the reference sample $\gamma_{0}$ into the data sample $\gamma_{1}$.
We now lift this deterministic path $(\gamma_{s}^{z})_{s=0}^{1}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s=0}^{1}\subset\mathbb{P}(\Gamma)$. We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each of the events in $\gamma_{s}^{z}$. That is, a draw $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ may be simulated via
@@
In principle this means that the support of $\eta_{s}^{z}$ is larger than $\mathcal{T}$, but in practice we choose $\sigma^{2}$ sufficiently small such that this is not a concern. The addition of noise $\epsilon_{k}$ is instrumental in obtaining a well-specified model, but in practice the noise variance $\sigma^{2}$ is not a critical hyperparameter. We note that this noising step is typical in flow matching models (Lipman et al., 2023; Tong et al., 2024).
@@
Finally, for $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}\in\mathbb{P}(\Gamma)$ via $\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$. Observe that, by construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$. This path of TPP distributions $\eta_{s}$ approximately interpolates from the reference TPP $\mu_{0}$ at $s=0$ to the data TPP $\mu_{1}$ at $s=1$, in the sense that at the endpoints, the joint event time distributions $\eta_{0}^{n}(t^{1},\dots,t^{n})$ and $\eta_{1}^{n}(t^{1},\dots,t^{n})$ are given by a convolution of $\mu_{0}^{n}(t^{1},\dots,t^{n})$ and $\mu_{1}^{n}(t^{1},\dots,t^{n})$ with the Gaussian $\mathcal{N}(0,\sigma^{2}I_{n})$. As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant (Tong et al., 2024).
For single event $t_{s}^{k}$, we have constructed a path of distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$. This transport of a Gaussian can be achieved infinitesimally through the vector field $v_{s}^{k}(t)=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024). Thus, the evolution (4) is generated by the vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$,
$v_{s}^{z}(\gamma)=\begin{bmatrix}v_{s}^{1},\dots,v_{s}^{n}\end{bmatrix}^{\sf T}=\begin{bmatrix}t_{1}^{1}-t_{0}^{1},&\dots,&t_{1}^{n}-t_{0}^{n}\end{bmatrix}^{\sf T}\qquad 0\leq s\leq 1.$ (5)
In other words, flowing samples $\hat{\gamma}_{0}^{z}$ along the vector field $v_{s}^{z}$ generates the path of distributions $\eta_{s}^{z}$. However, this path is conditioned on $z$, and we would like to find the vector field $v_{s}$ which generates the unconditional path $\eta_{s}$. As is standard in flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023), the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ may be obtained via
$v_{s}(\gamma)=\int v_{s}^{z}(\gamma)\frac{\,\mathrm{d}\eta_{s}^{z}}{\,\mathrm{d}\eta_{s}}(\gamma)\,\mathrm{d}\rho(z).$ (6)
We have thus far described a procedure for interpolating between a given reference distribution $\mu_{0}^{n}$ and the data distribution $\mu_{1}^{n}$ for a given, fixed number of events $n$. As $n$ was arbitrary, we have successfully constructed a family of interpolants which will enable us to sample from the joint event distribution for any $n$. However, to fully characterize the TPP distribution, we need to also specify the event count distribution. For unconditional generation tasks, this is straightforward – we simply follow the empirical event count distribution see in the training data. We describe our approach for modeling the event count distribution in conditional tasks in the following section.
1 Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, $\epsilon\sim\mathcal{N}(0,1)$
2 $e_{\mathcal{H}}=\varnothing$
3  /* Null history */
4 if forecast then
5       Sample $T_{0}\in[\Delta T,T-\Delta T]$
6       $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
7       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
8       $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
9
[…]
11 Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
13 $t_{s}^{k}\leftarrow(1-s)t_{0}^{k}+st_{1}^{k}$ for $k=1,\dots,n$
14 $\gamma_{s}^{z}\leftarrow(t_{s}^{1},\dots,t_{s}^{n})$
@@
1 Choose a flow time discretization $10=s_{0}<s_{1}<\dots<s_{K}=$ 1
2 $e_{\mathcal{H}}=\varnothing$
3  /* Null history */
4 if forecast then
5       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H}$)
6       $n\sim p_{\phi}(n\mid\mathcal{H})$
7
8else
9       $n\sim\mu_{1}(n)$
10
11Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
13 for $k=1,2,\dots,K$ do
14       $h_{k}\leftarrow s_{k}-s_{k-1}$
15       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
16
@@ 15       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
Training
If we knew $v_{s}$ in Equation (6), we could draw samples from $\mu_{1}$ by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP and flowing each event along $v_{s}$. Although the marginal vector field in Equation (6) admits an analytical form, it is intractable to compute in practice as the marginal measure $\eta_{s}$ is not available. To overcome this, we instead regress on the conditional vector fields $v_{s}^{z}$. Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ along with the flow time $s$. We seek to minimize the loss
which previous works on flow matching have shown to be equal to MSE regression on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024). We estimate $J(\theta)$ by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced coupling and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (4).
To train the model on a forecasting task, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (7). Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$. In this case, we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$. We train $p_{\phi}(n\mid\mathcal{H})$ by minimizing a cross-entropy loss. In practice, we found it important to regularize this loss to encourage smoothness in $n$, which we achieve through an optimal-transport based regularizer. As both $v_{\theta}$ and $p_{\phi}$ must be able to operate on variable-length sequences, we use a transformer-based backbone for both models. However, our overall method is agnostic to this choice. See Algorithm 1 and Appendix C.

### v3

@@ 4 EventFlow
Motivated by the limitations of autoregressive models, we propose EventFlow, which has a number of distinct advantages over prior work. First, EventFlow directly models the joint distribution over event times, thereby avoiding autoregression entirely. Second, our model is likelihood-free at training time, avoiding the Monte Carlo estimates needed to estimate the likelihood in Equation (3). Third, sampling from our model amounts to solving an ordinary differential equation. This is straightforward to implement and parallelize, allowing us to avoid the difficulties of thinning-based approaches. In fact, EventFlow can achieve competitive performance using only a single forward pass at sampling time.
We build upon the flow matching (or stochastic interpolant) framework (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) to develop our model. We begin below by focusing on the unconditional setting (i.e., generating an event sequence without being given an observed history), followed by an extension for conditional generation.
@@
We first introduce some necessary background and notation. Let $\mathcal{T}=[0,T]$ be a finite length interval. The set $\Gamma$ denotes the configuration space of $\mathcal{T}$ (Albeverio et al., 1998), i.e., the set of all finite counting measures on the set $[0,T]$. A point $\gamma\in\Gamma$ corresponds to a measure of the form $\gamma=\sum_{k=1}^{n}\delta[t^{k}]$, i.e., a finite collection of Dirac deltas located at event times $t^{k}\in\mathcal{T}$. A temporal point process (TPP) on $\mathcal{T}$ is a probability distribution $\mu$ over the configuration space $\Gamma$. We use $\mathbb{P}(\Gamma)$ to denote the set of all such distributions. Informally, a TPP $\mu$ is a distribution over finite sequences $\gamma\in\Gamma$ whose events are in $\mathcal{T}$. We use $N:\Gamma\to\mathbb{Z}_{\geq 0}$ to denote the counting functional, i.e., $N(\gamma)$ is the number of events in the sequence $\gamma$. [^1] This can be thought of in terms of the counting process, i.e., $N(\gamma)$ corresponds to the value of the associated counting process at the ending time $T$, or the total number of events in $\gamma$ that have occurred in the interval $[0,T]$. While it is common to represent TPPs as distributions over random sets of event times, in our approach it will be more convenient to represent TPPs as random measures (Kallenberg, 2017).
Under mild assumptions, a TPP $\mu$ can be fully characterized (Daley and Vere-Jones, 2003, Prop. 5.3.II) by a probability distribution which specifies the number of events and a collection of joint densities corresponding to the event times themselves. In a slight abuse of notation, we will write $\mu(n)$ for the corresponding distribution over event counts, and $\{\mu^{n}(t^{1},\dots,t^{n})\}_{n=1}^{\infty}$ for the collection of joint distributions. In other words, for any given $n\in\mathbb{Z}_{\geq 0}$, the probability of observing $n$ events in the interval $\mathcal{T}$ is $\mu(n)$, and $\mu^{n}(t^{1},\dots,t^{n})$ describes the corresponding joint distribution of event times. We further restrict each $\mu^{n}$ to be supported only on the ordered sets, so that $t^{1}<t^{2}<\dots<t^{n}$.
Let $\mu_{1}$ represent the data distribution and $\mu_{0}$ represent a reference distribution, i.e., both $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ are distributions over event sequences. To construct our model, we will define a path of distributions $(\eta_{s})_{s\in[0.1]}\subset\mathbb{P}(\Gamma)$ which approximately interpolates from our reference TPP to our data TPP. Intuitively, this path provides us with a way to transform samples: we can draw a sequence of events from $\mu_{0}$ and gradually move them along this path to obtain a sequence approximately distributed according to $\mu_{1}$. Throughout, we use $s\in[0,1]$ to denote a flow time and $t\in[0,T]$ to denote an event time. These two time axes are in a sense orthogonal to one another (see Figure 1).
@@
Towards constructing such a path, our first step is to define a notion of coupling, which allows us to pair event sequences drawn from the reference distribution $\mu_{0}$ with those drawn from the data distribution $\mu_{1}$. Formally, a coupling between two TPPs $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ is a joint probability distribution $\rho\in\mathbb{P}(\Gamma\times\Gamma)$ over pairs of event sequences $(\gamma_{0},\gamma_{1})$ such that the marginal distributions of $\rho$ are $\mu_{0}$ and $\mu_{1}$. Sampling $(\gamma_{0},\gamma_{1})\sim\rho$ from a coupling gives us paired sequences of events.
While couplings are broadly used in transport-based generative modeling (Tong et al., 2024; Villani et al., 2009), we introduce a new class of couplings which is particularly well-suited to the TPP setting. We say that the coupling $\rho$ is balanced if draws $(\gamma_{0},\gamma_{1})\sim\rho$ are such that $N(\gamma_{0})=N(\gamma_{1})$ almost surely. That is, balanced couplings only pair event sequences with equal numbers of events. This ensures that when we later construct interpolations between paired sequences, we do not need to add or remove events, simplifying both training and sampling. We will use $\Pi_{b}(\mu_{0},\mu_{1})$ to denote the set of balanced couplings.
The following proposition characterizes when balanced couplings exist. In particular, $\Pi_{b}(\mu_{0},\mu_{1})$ is nonempty if and only if the event count distributions of $\mu_{0}$ and $\mu_{1}$ are identical, imposing a structural constraint on suitable choices of $\mu_{0}$: the reference TPP $\mu_{0}$ must have the same event count distribution as the data TPP $\mu_{1}$.
@@
Let $\mu_{0},\mu_{1}\in\mathbb{P}(\Gamma)$ be two TPPs. The set of balanced couplings $\Pi_{b}(\mu_{0},\mu_{1})$ is nonempty if and only if have the same distribution over event counts, i.e., $\mu_{0}(n)=\mu_{1}(n)$ for every $n\in\mathbb{Z}_{\geq 0}$.
We provide a proof in Appendix B.
In practice, we follow a simple strategy for choosing both the reference TPP $\mu_{0}$ and the coupling $\rho$. Let $q$ be a density on $\mathcal{T}$ (e.g., uniform). We take $\mu_{0}$ to be a mixed binomial process (Kallenberg, 2017, Ch. 3) whose event count distribution is given by that of the data $\mu_{1}(n)$, and joint event time distributions given by independent products of $q$ (up to sorting). Sampling $\gamma_{0}\sim\mu_{0}$ amounts to sampling $n\sim\mu_{1}(n)$ from the data event count distribution followed by sampling and sorting $n$ i.i.d. times $t^{k}\sim q$. Under this choice, a sample from a balanced coupling $\rho\in\Pi_{b}(\mu_{0},\mu_{1})$ can be produced by first sampling a data sequence $\gamma_{1}\sim\mu_{1}$, followed by sampling $N(\gamma_{1})$ events independently from $q$ and sorting to produce a paired draw $\gamma_{0}\sim\mu_{0}$.
@@
We now construct our path $(\eta_{s})_{s\in[0,1]}\subset\mathbb{P}(\Gamma)$ using a two-stage procedure. First, for a given pair of sequences $(\gamma_{0},\gamma_{1})$ drawn from a balanced coupling, we define a sequence-level interpolation $(\gamma_{s})_{s\in[0,1]}$ that smoothly transforms $\gamma_{0}$ into $\gamma_{1}$. Second, the distribution $\eta_{s}$ is obtained by averaging these interpolated sequences over the coupling. We extend flow matching techniques (Lipman et al., 2023; Tong et al., 2024) to define the interpolants, and emphasize that fixing the number of events via a balanced coupling is essential for this construction.
To that end, let $\rho$ be any balanced coupling of the reference distribution $\mu_{0}$ and the data distribution $\mu_{1}$, and suppose $z:=(\gamma_{0},\gamma_{1})\sim\rho$ is a draw from this coupling. As $\rho$ is balanced, we have $\gamma_{0}=\sum_{k=1}^{n}\delta[t_{0}^{k}]$ and $\gamma_{1}=\sum_{k=1}^{n}\delta[t_{1}^{k}]$ are both a collection of $n$ events. We will henceforth describe our procedure for a fixed (but arbitrary) number of events $n$, and we will later describe how to model the number of events itself. First, we define the interpolant sequence $\gamma_{s}^{z}\in\Gamma$ via
@@
where we use the superscript $z$ to denote the dependence on the pair $z=(\gamma_{0},\gamma_{1})$. In other words, $\gamma_{s}^{z}$ linearly interpolates each corresponding event in $\gamma_{0}$ and $\gamma_{1}$, defining a path $(\gamma_{s}^{z})_{s\in[0,1]}\subset\Gamma$ which evolves the reference sample $\gamma_{0}$ into the data sample $\gamma_{1}$.
We now lift this deterministic path $(\gamma_{s}^{z})_{s\in[0,1]}\subset\Gamma$ to a path of TPP distributions $(\eta_{s}^{z})_{s\in[0,1]}\subset\mathbb{P}(\Gamma)$. We define the point process distribution $\eta_{s}^{z}\in\mathbb{P}(\Gamma)$ implicitly by adding independent Gaussian noise to each event in $\gamma_{s}^{z}$. That is, a draw $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ may be simulated via
@@
In principle, using Gaussian noise means that the support of event times in $\eta_{s}^{z}$ is larger than $\mathcal{T}$, but in practice we choose $\sigma^{2}$ sufficiently small such that this is not a concern. The addition of noise $\epsilon^{k}$ is instrumental in obtaining a well-specified model, but in practice we found the noise variance $\sigma^{2}$ to not be a critical hyperparameter. We note that this noising step is typical in flow matching models (Lipman et al., 2023; Tong et al., 2024).
Finally, for $s\in[0,1]$, we define the marginal TPP measure $\eta_{s}\in\mathbb{P}(\Gamma)$ via the mixture distribution
$\eta_{s}=\int\eta_{s}^{z}\,\mathrm{d}\rho(z)$ (7)
which, intuitively, corresponds to the two-stage procedure of drawing a sample $z\sim\rho$ and subsequently following the conditional flow defined in Equation (6).
@@
By construction, the event count distribution $\eta_{s}(n)$ is given by $\mu_{1}(n)$ for all $s\in[0,1]$. This path of TPP distributions $\eta_{s}$ approximately interpolates from the reference TPP $\mu_{0}$ at $s=0$ to the data TPP $\mu_{1}$ at $s=1$, in the sense that at the endpoints, the joint event time distributions $\eta_{0}^{n}(t^{1},\dots,t^{n})$ and $\eta_{1}^{n}(t^{1},\dots,t^{n})$ are given by a convolution of $\mu_{0}^{n}(t^{1},\dots,t^{n})$ and $\mu_{1}^{n}(t^{1},\dots,t^{n})$ with the Gaussian $\mathcal{N}(0,\sigma^{2}I_{n})$. As $\sigma^{2}\downarrow 0$, it is clear that we recover a genuine interpolant.
Observe that in Equation (6), for each individual event $t_{s}^{k}$, we have constructed a path of Gaussian distributions $\mathcal{N}(t_{s}^{k},\sigma^{2})$ centered around said event. This path of Gaussians can be simulated via the vector field $v_{s}^{k}:=t_{1}^{k}-t_{0}^{k}$ (Tong et al., 2024). That is, if we draw an initial condition $\tau_{0}^{k}\sim\mathcal{N}(t_{0}^{k},\sigma^{2})$ and solve the differential equation $\mathrm{d}\tau_{s}^{k}=v_{s}^{k}\,\mathrm{d}s$ for $s\in[0,1]$ starting from $\tau_{0}^{k}$, we obtain a draw $\tau_{1}^{k}\sim\mathcal{N}(t_{1}^{k},\sigma^{2})$. Loosely, $v_{s}^{k}$ is a velocity associated with the event $t_{s}^{k}$.
Since each event in Equation (6) evolves independently, the path of distributions ${\eta}_{s}^{z}$ is generated by the constant vector field $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ defined as
$v_{s}^{z}(\gamma):=\begin{bmatrix}v_{s}^{1},\dots,v_{s}^{n}\end{bmatrix}^{\sf T}=\begin{bmatrix}t_{1}^{1}-t_{0}^{1},&\dots,&t_{1}^{n}-t_{0}^{n}\end{bmatrix}^{\sf T}.$ (8)
In other words, flowing initial samples $\hat{\gamma}_{0}^{z}\sim\eta_{0}^{z}$ along the vector field $v_{s}^{z}(\gamma)$ generates the path of distributions $\eta_{s}^{z}$, approximately interpolating from the noise sequence $\gamma_{0}$ to the data sequence $\gamma_{1}$.
However, this path is conditioned on $z=(\gamma_{0},\gamma_{1})$. Since we seek to generate $\gamma_{1}$, this is intractable, and we would instead like to find the vector field $v_{s}$ that generates the unconditional path $\eta_{s}$ defined in Equation (6). To this end, we aggregate the conditional vector fields $v_{s}^{z}(\gamma)$ across the coupling $z\sim\rho$. In particular, the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ is given by
$v_{s}(\gamma)=\mathbb{E}_{z\sim\rho}[v_{s}^{z}(\gamma)\mid\gamma_{s}^{z}=\gamma]=\int v_{s}^{z}(\gamma)\frac{\,\mathrm{d}\eta_{s}^{z}}{\,\mathrm{d}\eta_{s}}(\gamma)\,\mathrm{d}\rho(z)$ (9)
i.e., the average of the conditional vector fields over all pairs $z$ that could have produced the sequence $\gamma$ at time $s$. This corresponds to an extension of the standard flow matching construction (Lipman et al., 2023; Tong et al., 2024; Albergo and Vanden-Eijnden, 2023) to the setting of TPPs.
Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, $\epsilon\sim\mathcal{N}(0,1)$
$e_{\mathcal{H}}=\varnothing$
/* Null history */
if forecast then
   Sample $T_{0}\in[\Delta T,T-\Delta T]$
 $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
 $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
 $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
end if
$n\leftarrow N(\gamma_{1})$
Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
$\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
$t_{s}^{k}\leftarrow(1-s)t_{0}^{k}+st_{1}^{k}$ for $k=1,\dots,n$
$\gamma_{s}^{z}\leftarrow(t_{s}^{1},\dots,t_{s}^{n})$
@@
Training
If we knew $v_{s}$ in Equation (9), we could draw a sample from $\gamma_{1}\sim\mu_{1}$ by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP and flowing each event along $v_{s}$. More precisely, $\gamma_{0}$ will consist of $N(\gamma_{0})$ events and $v_{s}(\gamma)$ will be a vector field with $N(\gamma_{0})$ components, so that we may solve the differential equation $\mathrm{d}\gamma_{s}=v_{s}(\gamma)\,\mathrm{d}s$ on $s\in[0,1]$.
Although the marginal vector field in Equation (9) admits an analytical form, it is intractable to compute in practice as the density ratio $\mathrm{d}\eta_{s}^{z}/\,\mathrm{d}\eta_{s}$ is unknown. To overcome this, we instead regress on the conditional vector fields $v_{s}^{z}$. Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ along with the flow time $s$. We aim to minimize
which is equal to the MSE loss on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024). We estimate $J(\theta)$ by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced coupling and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (6).
To train the model for forecasting, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (10). Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$. See Algorithm 1 for training pseudocode.
Event Count Distributions
We have thus far described a procedure for interpolating between a given reference distribution $\mu_{0}^{n}$ and the data distribution $\mu_{1}^{n}$ for a given, fixed number of events $n$. As $n$ was arbitrary, we have successfully constructed a family of interpolants which will enable us to sample from the joint event time distributions $\mu_{1}^{n}(t^{1},\dots,t^{n})$. However, recall that fully characterizing the TPP distribution requires us to also specify the event count distribution.
For unconditional generation tasks, this is straightforward: we simply follow the empirical event count distribution see in the training data. When forecasting, though, we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$, i.e., how many events are likely to occur after observing a history $\mathcal{H}$. Here, $\phi$ represents the parameters of this model. We train $p_{\phi}(n\mid\mathcal{H})$ by minimizing a cross-entropy loss. In practice, we found it important to regularize this loss to encourage smoothness in $n$, which we achieve through an optimal-transport based regularizer. See Appendix C for details.
As both $v_{\theta}$ and $p_{\phi}$ must be able to operate on variable-length sequences, we use a transformer-based backbone for both models. However, our overall method is agnostic to this choice.
Choose a flow time discretization $0=s_{0}<s_{1}<\dots<s_{K}=1$
$e_{\mathcal{H}}=\varnothing$
/* Null history */
if forecast then
 $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
 $n\sim p_{\phi}(n\mid\mathcal{H})$
else
 $n\sim\mu_{1}(n)$
end if
Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
$\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
for $k=1,2,\dots,K$ do
 $h_{k}\leftarrow s_{k}-s_{k-1}$
 $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
end for